OpenStudy (anonymous):

providing the formula would be helpful. John wishes to set up an account for his grandfather so that he can have some extra money each month. John wants his grandfather to be able to withdraw $140 per month for the next 4 years. How much must John invest today at 8% per year compounded monthly so that his grandfather can withdraw $140 per month for the next 4 years? a) $5,714.67 b) $5,764.67 c) $5,694.67 d) $5,734.67 e) $5,724.67

OpenStudy (anonymous):

Well, for compounding interest, we have the formula: \[ I = Pr^t \]Where \(I\) is the interest made, \(P\) is the principle (initial amount of money), \(r\) is the interest rate, and \(t\) is the number of times it's compounding. Thought this problem will require a bit more thinking.

OpenStudy (anonymous):

If you want the new ammount, it's just interest plug principle.\[ P+I = P(1+r)^t \]

OpenStudy (mathstudent55):